OpenStudy (anonymous):

Evaluate the function rule for the given value. f(x) = 4^x for x = 2. -16 16 4 8 I came up with 16 am I right or wrong on this one?

OpenStudy (zzr0ck3r):

\(f(2) = 4^2 = 16 \checkmark\)

OpenStudy (zzr0ck3r):

yep

OpenStudy (anonymous):

\[f(x) = 4^x~~~x = 2~~~~f(2) = 4^2 = 4 \times 4 = 16\]

OpenStudy (anonymous):

thankyou everyone :)